Compute the total and annual return on the following investment.Five years after paying

​$3000 for shares in a startup​ company, you sell the shares for $2000 (at a​ loss).

The total return is _%.?

​(Do not round until the final answer. Then round to the nearest tenth as​ needed.)
Answer #1

a)Total return = [P1-P0]/P0

                 = [2000-3000]/3000

                   = -1000/3000

                  = - .333 or -33.3%

b)

Future value =Present value (1+i)^n

2000 = 3000(1+i)^5

2000/3000 = (1+i)^5

(.66667) = (1+i)^5

multiplying both sides by ^1/5

(.66667)^1/5 = (1+i)^1/5*5

(.66667)^.20 = (1+i)

.9221 = (1+i)

i = .9221 -1

   = -.0779 or -7.79% (rounded to -7.8%)

annual return = -7.8%